OpenStudy (anonymous):

help if possible? Write x2 - 2x - 3 = 0 in the form (x - a)2 = b, where a and b are integers. (4 points) (x - 4)2 = 3 (x - 3)2 = 2 (x - 2)2 = 1 (x - 1)2 = 4

OpenStudy (decentnabeel):

\[\left(x-1\right)^2=4\]

OpenStudy (decentnabeel):

subtract 3 on both sides x^2 - 2x = -3 \[\mathrm{Add\:}\left(\frac{-2}{2}\right)^2\mathrm{\:\to\:both\:sides}\] \[x^2-2x+1=4\] \[\left(x-1\right)^2=4\]

OpenStudy (anonymous):

thank you

OpenStudy (decentnabeel):

wellcome :) @kinserkara